OpenStudy (anonymous):

I need someone to check my answers, and maybe help me with a few more? Please? My questions are from FLVS.

OpenStudy (anonymous):

You and some friends have started your own company. After the first few months, the profits are rolling in. It is time to start thinking about putting your money to work for you. You decide that investing $5,000 into some Certificates of Deposit (CDs) would be a beneficial move. With a CD, you lend your money to a third party, and after a set time, your money is paid back with interest. Before you start investing the company's money this way, you need to pitch it to your friends. For this project, you can develop a written report, a slideshow, a video, or any other platform as long as it fully addresses the criteria listed below. 1. Research the highest interest rate (APY—annual percent yield) for 2-year and 5-year CDs. Document the company's name, interest rate, and minimum investment. The minimum investment must be less than or equal to $5,000. Answer: My research showed that the only bank that does 2-year and 5-year CDs that does not require a minimum deposit of $25,000.00 is Ally Bank. Ally Bank’s interest rate for a 2-year CD is 1.050% and for a 5-year CD it is 1.600%. The minimum investment is $0.00. 2. Create the functions that represent the 2-year and 5-year CDs with your $5,000 investment. Use these functions to determine the amount you will be paid when the CD matures (the length of time for the specific CD). Show your work. Answer: I made the 2-year CD f(x) and the 3-year CD y(x). f(x)=5000(1.050)+0 y(x)=5000(1.600)+0

OpenStudy (anonymous):

@timaashorty Do you think you can help? If these are correct, Awesome, but I cannot understand the other. Ill copy and paste them.